OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 5. Find the number.

OpenStudy (anonymous):

\[x+\frac{6}{x} = 5\] \[ \frac{x^{2}+6}{x} = 5\]

OpenStudy (anonymous):

x^2 + 6 =5x x^2 - 5x + 6 =0 (x-3)(x-2) = 0 x = 3, 2